# pacer: Preference-conditioned All-terrain Costmap Generation

### v1

@@ Abstract
In autonomous robot navigation, terrain cost assignment is typically performed using a semantics-based paradigm in which terrain is first labeled using a pre-trained semantic classifier and costs are then assigned according to a user-defined mapping between label and cost. While this approach is rapidly adaptable to changing user preferences, only preferences over the types of terrain that are already known by the semantic classifier can be expressed. In this paper, we hypothesize that a machine-learning-based alternative to the semantics-based paradigm above will allow for rapid cost assignment adaptation to preferences expressed over new terrains at deployment time without the need for additional training. To investigate this hypothesis, we introduce and study pacer, a novel approach to costmap generation that accepts as input a single birds-eye view (BEV) image of the surrounding area along with a user-specified preference context and generates a corresponding BEV costmap that aligns with the preference context. Using both real and synthetic data along with a combination of proposed training tasks, we find that pacer is able to adapt quickly to new user preferences while also exhibiting better generalization to novel terrains compared to both semantics-based and representation-learning approaches.
Figure 1: Given an input image $I$ and a preference context $\hat{H{}}$ of $n$ ordered pairs of terrain patches where the left terrain is more preferred than the right, pacer generates a costmap consistent with this preference. Changing the preference context leads to changed terrain costs, which results in a different plan aligned to the new operator preference. The paths planned according to the different preferences are shown above.
@@ I Introduction and Related Work
Robust and aligned autonomous ground vehicle navigation in a wide variety of environments is a long-standing goal in the robotics community. While there has been progress on the classical problem of collision-free waypoint navigation in static, simple environments [1][2], there are still significant challenges in constructing autonomous systems that can navigate in environments that are dynamic, more complex, or both [3]. Further, successful operation in such environments typically relies on robots understanding and aligning their behaviors to specified human preferences for navigation that go beyond simple waypoint achievement, e.g., preferring to cross a busy street at a crosswalk even if doing so results in a longer path to a waypoint [4][5].
In this paper, we consider the specific question of how robots should assign costs to terrain such that those costs align with human preferences for terrain traversal. Having robots adhere to preferences of this type is useful both for helping enable robust navigation in complex environments and also for ensuring that robots adhere to other constraints such as social norms [6] [7]. While seeking to assign appropriate costs to terrain is not the only way to ensure aligned navigation behavior (e.g., one may instead seek navigation control policies that are aligned to human preferences [8][9][10]), we choose to focus on assigning terrain costs here due to its prevalence in the existing literature, as many existing and well-understood autonomous planning frameworks leverage costmaps to produce cost-optimal behaviors [11] [12] [13].
We are particularly interested here in terrain cost assignment approaches that can rapidly adapt to newly-expressed terrain preferences. Popular approaches such as inverse reinforcement learning (IRL) and preference-based IRL (PbIRL) based on terrain patch clusters do not admit this type of rapid adaptation due to the amount of additional data required to express or query for the new preference [14] [15] [16]. Instead, to the best of our knowledge, the most commonly-employed systems with this capability follow a semantics-based paradigm in which the surrounding terrain is first labeled using a pre-trained semantic classifier and human preferences are conveyed via a manually specified label-to-cost mapping [5] [17][18]. Such systems exhibit rapid adaptation to new preferences by design: the new costmaps can be generated as soon as the user provides a new label-to-cost mapping, which is typically accomplished in seconds to minutes. However, these systems are also inherently limited in that the user can only specify preferences with respect to the terrain types known to the classifier.
Another class of methods designed for terrain-aware aware navigation is by representing terrains as vectors in a continuous embedding space, so preference are no no longer limited to terrains with predefined labels [14] [19] [20]. Patch-based representation learning methods for terrain understanding typically involve mapping small square patches from the bird’s-eye view (BEV) to a representation vector that is further converted into a scalar cost value.
@@
Although continuous representation spaces are theoretically generalizable to new terrains, training such a space effectively is challenging in practice, since even humans may struggle to identify terrain types from small patches in the presence of homography artifacts or difficult lighting conditions. A further limitation is that each new preference ordering over terrains necessitates retraining the utility function in the representation space, making these approaches less adaptable to changing operator preferences.
@@ II-B Preference-Aligned Terrain Costs
For terrain-aware navigation, the robot relies on its visual observations to infer terrain-specific costs. We assume that these observations arrive in the form of images generated according to a black-box observation function $O{}:\mathcal{X}{}\times\mathbb{T}{}\to\mathcal{I}{}$, i.e., $I=O(x,T)$, where $x$ is the observing pose of the robot, $T$ is a terrain map, $\mathbb{T}$ is the space of terrain maps, and $\mathcal{I}$ is the space of images. In practice, most methods operate on synthetic birds-eye-views generated from the original camera images. BEV images can be generated via static ground-plane homography [15][14], or a BEV accumulation algorithm [21]. Henceforth, we define input images to be BEV images. We assume that the the visual appearance of the terrain provides sufficient information for the robot to perform terrain-aware navigation. The observation function is thus fixed, but unknown to the robot.
@@ III Necessary Conditions for Preference-Aligned Navigation
Seeking training tasks that will help us compute valid preference-conditioned costmap functions $R(\cdot|H)$, we now state a set of conditions that are necessary for these tasks to produce costmaps that are consistent with human preferences for terrain.
In particular, we will state conditions for equivalence, partial ordering, and viewpoint invariance, and we will show that $R$s that produce costmaps that yield optimal trajectories consistent with a human preference must obey these conditions.
Let $R{}(\cdot|H{})$ denote a costmap generated according to an $H{}\in\mathcal{H}{}$, and $C{}|_{x}{}$ denote that costmap $C{}\in\mathcal{C}{}$ is evaluated at pose $x$. For a generated costmap $R{}(\cdot|H{})$ to be consistent with $H{}$
, we specify it must exhibit both equivalence and partial ordering. By equivalence, we mean that the terrains at two poses are given the same cost by $H$ if and only if the costmap generated by $R$ from an image observation and evaluated at those two poses have equal cost, i.e.,
@@
We now provide a brief proof that NC1 NC2 are necessary for aligning the preferences of $H{}_{i}$ with $R{}$. Specifically, if the most optimal path with respect to $R{}(\cdot\mid H{}_{i})$ has the same optimal cost when evaluated with $H{}_{i}$, then the conditions (NC1) and (NC2) must hold. For a trajectory $\Gamma$ composed of discrete poses, let the cumulative cost function for the human’s evaluation be denoted by $H{}|_{\Gamma}\equiv\sum_{x_{i}\in\Gamma}H{}(T(x{}_{i}))$. Similarly, let the cumulative cost function based on the generated costmap be denoted as $R{}|_{\Gamma}\equiv\sum_{x_{i}\in\Gamma}R{}(O{}(\cdot,T)\mid H{})|_{x_{i}}$. Given this setup, the following theorem establishes the necessity of the conditions (NC1) and (NC2) such that $H{}|_{\Gamma^{\ast}}=R{}|_{\overline{\Gamma}}$.
@@ IV-A Preference Context
such that the human prefers the terrain observed in image $\widetilde{I{}}$ over the terrain observed in $\widetilde{I{}}^{\prime}$, where an $\widetilde{I{}}\in\mathcal{I}{}$ is an observation of terrain as a small image patch. The small patch may be a part of a larger bird’s-eye-view image of the ground.
More specifically, $\hat{H{}}$ consists of $n$ preferences derived from $H{}$ and is defined as $\hat{H{}}\equiv\{(\widetilde{I{}}_{1}\succ\widetilde{I{}}_{1}^{\prime}),...(\widetilde{I{}}_{n}\succ\widetilde{I{}}_{n}^{\prime})\}$. Fig 3 shows some example preference contexts with $n=3$ patch pairs and their corresponding costmaps.
In our implementation, the $n$ pairwise preferences are expressed using image patches of size $w\times h$. $\hat{H{}}$ is then represented by vertically concatenating the patches within a pair with the more-preferred terrain patch on top and forming a single $n\times c\times w\times 2h$ tensor, where $c$ is the number of color channels.
@@ V Dataset Curation and Training pacer
We now describe the dataset curation and training process for pacer. pacer is trained using three distinct phases of supervised machine learning, each corresponding to a unique training dataset that corresponds to one of the desiderata desecribed in Section I. In what follows, we will first describe how we generate training examples, then describe each of the three training phases and the training procedure.
@@ V-A Training Example Generation
The inputs to the training example generation process are a single image $I$ along with a total ordering over terrain types $\tau_{1}\succ\tau_{2}\succ\ldots\succ\tau_{L}$, where each terrain type $\tau_{l}$ corresponds to a bank of image patches. pacer assumes that the cost value associated with $\tau_{l}$ is given by $H{}(\tau_{l})=\frac{l-1}{L-1}$. The patch bank that corresponds to $\tau_{l}$ consists of patches of that type that have been extracted from images collected during robot deployment.
We use these inputs to generate $\hat{H{}}$ and $C{}_{T}$. To generate $\hat{H{}}$, we first choose $n$ ordered pairs from the total ordering over the $L$ terrain types without replacement. For each of the resulting ordered pairs, we sample uniformly at random patches from the corresponding patch banks, and use these $2n$ patches to construct $\hat{H{}}$ according to the process detailed in Section IV-A above. To generate $C{}_{T}$, we first perform semantic segmentation on $I$ and then transform the segmented image into $C_{T}$ by setting the cost for a pixel labelled $l$ to be $H{}(\tau_{l})$.
@@ V-B Dataset Size
The size of the space from which we sample data is very large. From a total ordering of $L$ labels, there are $m=\binom{L}{2}$ different ordered pairs of terrains and $\binom{m}{n}$ different sets of $n$ pairs. For each set of $n$ pairs, there $n!$ ways to shuffle the pairs to construct the preference context, yielding $\binom{m}{n}\cdot n!$ possible preference contexts. Moreover, for each label in the preference context, we sample a patch from the patch bank. Our dataset contains a bank of around $800$ patches for each terrain label, and about $950$ full images. Therefore, for each total ordering, we have $\binom{\binom{L}{2}}{n}n!$ arrangements of labels into preference contexts, where we sample a patch from a bank of $800$ patches for each label. For $L$ labels, $n=L\log L$ pairs are needed to fully describe a total ordering, though we evaluate on a smaller $n=3$ due to size considerations for the model and dataset.
@@ V-C Training Phases
Training Phase 1: Pretraining with Real Data and Realistic Preferences. To promote a prior towards an overall “realistic” ordering (as per our first desired property), pacer’s first training phase constructs and utilizes a dataset $\mathcal{D}_{\mathrm{real}}$ generated using real-world data collected from robot deployments around our campus at The University of Texas at Austin and realistic preferences over terrain classes. An example of a realistic preference ordering is as follows: concrete $\succ$ pebble $\succ$ mulch $\succ$ grass $\succ$ marble $\succ$ bush. The “realistic preferences” were defined by the first author according to considerations for robot safety (e.g. preferring grass over loose marble rock for a wheeled robot) and societal norms (e.g. preferring concrete over grass to avoid trampling lawns, even though both terrains are relatively safe).
Training Phase 2: Augmentation with Changed Preferences. During deployment in terrains not seen during training, the robot should adhere to preferences given by the operator (as per our second desired property). Even when operator preferences contradict “realistic preferences”, the robot should follow operator preferences over learned priors. To encourage this adherence to the ordering in the preference context, we train using the same real data but with changed preferences on a smaller corpus of data. More specifically, in this phase, we generate data using a randomly-permuted total ordering over terrain labels.
@@
We pick a random subset of terrains to replace and randomly permute the preference order. An image containing at least one such terrain is selected, and those terrains are artificially replaced with terrain textures from an open-source database [22] using dense segmentation. The training example is formed with a preference context (where terrains have been replaced), the image, and a costmap with costs reassigned according to the new preference order.
Training iterations are split equally among the three datasets $\mathcal{D}_{\mathrm{real}}$, $\mathcal{D}_{\mathrm{shuffled}}$, and $\mathcal{D}_{\mathrm{synthetic}}$. Before training on $\mathcal{D}_{\mathrm{real}}$, weights are initialized randomly. After completing a training phase, we switch to the next dataset starting from the previous trained weights.
@@ VI Experiments
By dividing deployment scenarios into the four situations above, we will be able to understand the performance of pacer under the four ways to combine seen and unseen terrains in the preference context and environment. In 1b and 2b, note that the preference context does not provide information about the terrains appearing in the environment, so the robot must rely on learned priors about realistic cost assignment.
Evaluations are performed using simulated experiments on an aerial map. In a later section, we also provide results from real robot deployments. We compare against sterling [14] (a representation learning approach) and a classifier (a semantics-based approach) as baselines. The same model for pacer is utilized across all environments. No retraining or fine-tuning is done. Additionally, a single context is used for the duration of a simulated deployment in an environment (i.e. the context is not switched out midway through the path-planning).
To quantify the navigation performance in our experiments, we posit that factors such as the distance traversed or closeness to a human-defined trajectory do not matter as much as traveling on only the preferred terrains. We therefore assign terrain types a low, medium, and high cost, and report the proportion of the planned path which traverses each of these terrain types. Note that we have purposely chosen this metric to be different from the commonly-used Hausdorff distance between the planned trajectory and one defined by a human operator, which can vary greatly when there are multiple valid paths to reach the goal.
Figure 4: The effect of changing preference on path planning is shown on the left image of an aerial map. The blue path corresponds to the middle costmap and the red path corresponds to the right costmap. The generated costmaps reflect the preferences provided in the context. Black corresponds to low cost and white corresponds to high cost.

### v2

@@ Abstract
In autonomous robot navigation, terrain cost assignment is typically performed using a semantics-based paradigm in which terrain is first labeled using a pre-trained semantic classifier and costs are then assigned according to a user-defined mapping between label and cost. While this approach is rapidly adaptable to changing user preferences, only preferences over the types of terrain that are already known by the semantic classifier can be expressed. In this paper, we hypothesize that a machine-learning-based alternative to the semantics-based paradigm above will allow for rapid cost assignment adaptation to preferences expressed over new terrains at deployment time without the need for additional training. To investigate this hypothesis, we introduce and study pacer, a novel approach to costmap generation that accepts as input a single birds-eye view (BEV) image of the surrounding area along with a user-specified preference context and generates a corresponding BEV costmap that aligns with the preference context. Using a staged training procedure leveraging real and synthetic data, we find that pacer is able to adapt to new user preferences at deployment time while also exhibiting better generalization to novel terrains compared to both semantics-based and representation-learning approaches. We release our code and dataset at https://github.com/ut-amrl/PACER_RAL_2025.git
Index Terms:
Vision-based navigation, Deep Learning for Visual Perception
@@ I Introduction and Related Work
Robust autonomous navigation in a wide variety of environments is a long-standing goal in robotics. While there has been significant progress in collision-free navigation [1, 2], successful navigation in human environments additionally requires alignment with human preferences, e.g., preferring to cross a busy street at a crosswalk even if doing so results in a longer path [3, 4].
In this paper, we examine how robots can assign terrain costs that align with human preferences for terrain-aware navigation. This specific focus represents a special case of the broader challenge of human preference-aligned navigation [5]. Such alignment is crucial not only for terrain-aware navigation, but also for adherence to constraints like social norms [6, 7]. An alternative to learning to predict navigation costs is to directly learn navigation policies aligned with human preferences [8, 9, 10]. We choose to focus on predicting navigation costs due to the ease of integration with existing cost-based navigation planners [11, 12, 13]. While the focus of this work is on terrain-based cost evaluation, we recognize that a full navigation system would incorporate additional costs using established methods such as layered costmaps [14].
We are particularly interested in terrain cost assignment approaches that can rapidly adapt to newly-expressed terrain preferences. Prevalent approaches to incorporate human preference into navigation such as inverse reinforcement learning (IRL) and preference-based IRL (PbIRL) based on terrain patch clusters do not admit this type of rapid adaptation due to the amount of additional data required to express new preferences [15, 16, 17]. Instead, to the best of our knowledge, existing solutions for rapid adaptation to preferences rely on first segmenting terrains into a prescribed set of classes, and assigning each class a manually-specified cost [4, 18, 19]. While such approaches allows rapid adaptation to new preferences, they are restricted to expressing preferences over the pre-defined list of terrain classes known to the segmentation algorithm.
Figure 1: Given an input image $I$ and a preference context $\hat{H{}}$ of $n$ ordered pairs of terrain patches where the left terrain is more preferred than the right, pacer generates a costmap consistent with this preference. Changing the preference context leads to changed terrain costs, which results in a different plan aligned to the new operator preference. The paths planned according to the different preferences are shown above. In the costmap, black represents low cost and white represents high cost.
@@
Representation learning is a more recent approach to terrain-aware navigation that allows preferences to no longer be limited to terrains with predefined labels [15, 20, 21, 16]. Patch-based representation learning methods for terrain understanding typically involve mapping small square patches from the bird’s-eye view (BEV) to a representation vector that is further converted into a scalar cost value. Although continuous representation spaces are theoretically generalizable to new terrains, training such a space effectively is challenging in practice, as even humans may struggle to identify terrain types from small patches in the presence of homography artifacts or difficult lighting conditions. A further limitation is that each new terrain preference ordering necessitates retraining the utility function in the representation space, making these approaches less adaptable to changing preferences.
@@ II-B Preference-Aligned Terrain Costs
For terrain-aware navigation, the robot relies on its visual observations to infer terrain-specific costs. We assume that these observations arrive in the form of images generated according to a black-box observation function $O{}:\mathcal{X}{}\times\mathbb{T}{}\to\mathcal{I}{}$, i.e., $I=O(x,T)$, where $x$ is the observing pose of the robot, $T$ is a terrain map, $\mathbb{T}$ is the space of terrain maps, and $\mathcal{I}$ is the space of images. In practice, most methods operate on synthetic birds-eye-views generated from the original camera images. BEV images can be generated via static ground-plane homography [16, 15], or a BEV accumulation algorithm [22]. Henceforth, we define input images to be BEV images. We assume that the the visual appearance of the terrain provides sufficient information for the robot to perform terrain-aware navigation. The observation function is thus fixed, but unknown to the robot.
@@ III Necessary Conditions for Preference-Aligned Navigation
Seeking training tasks to help us compute valid preference-conditioned costmap functions $R(\cdot|H)$, we now state a set of necessary conditions for these tasks to produce costmaps that are consistent with human preferences for terrain.
In particular, we will state conditions for equivalence and partial ordering, and we will show that $R$s that produce costmaps that yield optimal trajectories consistent with a human preference must obey these conditions.
Let $R{}(\cdot|H{})$ denote a costmap generated according to an $H{}\in\mathcal{H}{}$, and $C{}|_{x}{}$ denote that costmap $C{}\in\mathcal{C}{}$ is evaluated at pose $x$. For a generated costmap $R{}(\cdot|H{})$ to be consistent with $H{}$, we specify it must exhibit both equivalence and partial ordering. By equivalence, we mean that the terrains at two poses are given the same cost by $H$ if and only if the costmap generated by $R$ from an image observation and evaluated at those two poses have equal cost, i.e.,
@@
We now provide a brief proof that (NC1), (NC2) are necessary for aligning the preferences of $H{}_{i}$ with $R{}$. Specifically, if the most optimal path with respect to $R{}(\cdot\mid H{}_{i})$ has the same optimal cost when evaluated with $H{}_{i}$, then the conditions (NC1),(NC2) must hold. For a trajectory $\Gamma$ composed of discrete poses, let the cumulative cost function for the human’s evaluation be denoted by $H{}|_{\Gamma}\equiv\sum_{x_{i}\in\Gamma}H{}(T(x{}_{i}))$. Similarly, let the cumulative cost function for the generated costmap be denoted as $R{}|_{\Gamma}\equiv\sum_{x_{i}\in\Gamma}R{}(O{}(\cdot,T)\mid H{})|_{x_{i}}$. Given this setup, the following theorem establishes the necessity of the conditions (NC1), (NC2) such that $H{}|_{\Gamma^{\ast}}=R{}|_{\overline{\Gamma}}$.
@@ IV-A Preference Context
such that the human prefers the terrain observed in image $\widetilde{I{}}$ over the terrain observed in $\widetilde{I{}}^{\prime}$, where an $\widetilde{I{}}\in\mathcal{I}{}$ is an observation of terrain as a small image patch. The small patch may be a part of a larger bird’s-eye-view image of the ground. More specifically, $\hat{H{}}$ consists of $n$ preferences derived from $H{}$ and is defined as $\hat{H{}}\equiv\{(\widetilde{I{}}_{1}\succ\widetilde{I{}}_{1}^{\prime}),...(\widetilde{I{}}_{n}\succ\widetilde{I{}}_{n}^{\prime})\}$. Fig 3 shows some example preference contexts with $n=3$ patch pairs and their corresponding costmaps.
In our implementation, the $n$ pairwise preferences are expressed using image patches of size $h\times w$. $\hat{H{}}$ is then represented by vertically concatenating the patches within a pair with the more-preferred terrain patch on top and forming a single $(n\cdot c)\times 2h\times w$ tensor, where $c$ is the number of color channels. Given a finite preference context, it is impossible to specify all pairwise preferences over the terrains, especially since the terrain set is continuous - hence it is not possible to specify a stronger sufficient condition that would guarantee that the algorithm generates human-aligned costs.
@@ V Dataset Curation and Training pacer
We now describe the dataset curation and training process for pacer. pacer is trained using three distinct phases of supervised machine learning, each corresponding to a unique training dataset that corresponds to one of the desiderata described in Section I. In what follows, we will first describe how we generate training examples, then describe each of the three training phases and the training procedure.
@@ V-A Training Example Generation
The inputs to the training example generation process are a single image $I$ along with a total ordering over terrain types $\tau_{1}\succ\tau_{2}\succ\ldots\succ\tau_{L}$, where each terrain type $\tau_{l}$ corresponds to a bank of image patches. pacer assumes that the cost value associated with $\tau_{l}$ is given by $H{}(\tau_{l})=\frac{l-1}{L-1}$. The bank corresponding to $\tau_{l}$ consists of patches of that type extracted from images collected during robot deployment.
We use these inputs to generate $\hat{H{}}$ and $C{}_{T}$. To generate $\hat{H{}}$, we first choose $n$ ordered pairs from the total ordering over the $L$ terrain types without replacement. For each of the resulting ordered pairs, we sample uniformly at random patches from the corresponding patch banks, and use these $2n$ patches to construct $\hat{H{}}$ according to the process detailed in Section IV-A above. To generate $C{}_{T}$, we perform semantic segmentation on $I$ and transform the segmented image into $C_{T}$ by setting the cost for a pixel labelled $l$ to be $H{}(\tau_{l})$.
@@ V-B Dataset Size
The size of the space from which we sample data is very large. From a total ordering of $L$ discrete terrain types, there are $m=\binom{L}{2}$ different ordered pairs of terrains and $\binom{m}{n}$ different sets of $n$ pairs. For each set of $n$ pairs, there $n!$ ways to shuffle the pairs to construct the preference context, yielding $\binom{m}{n}\cdot n!$ possible preference contexts. Moreover, for each terrain type in the preference context, we sample a patch from the bank. Our dataset contains a bank of around $800$ patches for each terrain label, and about $950$ full images. Therefore, for each total ordering, we have $\binom{\binom{L}{2}}{n}n!$ arrangements of labels into preference contexts, where we sample a patch from a bank of $800$ patches for each terrain type. For $L$ terrains, $n=L\log L$ pairs are needed to describe a total ordering, though we evaluate on a smaller $n=3$ pairs due to size considerations for the model and dataset.
@@ V-C Training Phases
Training Phase 1: Pretraining with Real Data and Realistic Preferences. To promote a prior towards an overall “realistic” ordering (as per our first desired property), pacer’s first training phase constructs and utilizes a dataset $\mathcal{D}_{\mathrm{real}}$ generated using real-world data collected from robot deployments around our campus at The University of Texas at Austin and realistic preferences over terrain classes. An example of a realistic preference ordering is as follows: concrete $\succ$ pebble $\succ$ grass $\succ$ marble $\succ$ bush. The “realistic preferences” were defined by the first author according to considerations for robot safety (e.g. preferring grass over loose marble for a wheeled robot) and societal norms (e.g. preferring concrete over grass to avoid trampling lawns, even though both terrains are relatively safe).
Training Phase 2: Augmentation with Changed Preferences. During deployment in terrains not seen during training, the robot should adhere to preferences given by the operator (as per our second desired property). Even when operator preferences contradict “realistic preferences”, the robot should follow operator preferences over learned priors. To encourage this adherence to the ordering in the preference context, we train using the same real data but with changed preferences on a smaller corpus of data by using a randomly-permuted total ordering over terrain labels.
@@
We pick a random subset of terrains to replace and randomly permute the preference order. An image containing at least one such terrain is selected, and those terrains are artificially replaced with terrain textures from an open-source database [23] using dense segmentation. We used 14 synthetic textures. The training example is formed with a preference context (where terrains have been replaced), the image, and a costmap with costs reassigned according to the new preference order.
In the first phase, only $\mathcal{D}_{\mathrm{real}}$ is used for training examples. In the second, both $\mathcal{D}_{\mathrm{real}}$ and $\mathcal{D}_{\mathrm{shuffled}}$ are used. In the third, all three datasets are used. Within a phase, training examples are drawn uniformly among the datasets used. Before training on $\mathcal{D}_{\mathrm{real}}$, weights are initialized randomly. After completing a training phase, we switch to the next phase starting from the previous trained weights. We trained for 100 epochs in phase 1, 5 in phase 2, and 100 in phase 3.
@@ VI Experiments
By dividing deployment scenarios into the four situations above, we will be able to understand the performance of pacer under the four ways to combine seen and unseen terrains in the preference context and environment. In 1b and 2b, the preference context does not provide information about the terrains appearing in the environment, so the robot must rely on learned priors about realistic cost assignment. We term this scenario as having an uninformative context.
Evaluations are performed using simulated experiments on an aerial map. In a later section, we also provide results from real robot deployments. We compare against sterling [15] (a representation learning approach) and a classifier (a semantics-based approach) as baselines. The classifier is the same as used generate the training data for pacer. The same model for pacer is utilized across all environments. No retraining or fine-tuning is done. Additionally, a single context is used for the duration of a simulated deployment in an environment (i.e. the context is not switched out midway through the path-planning).
To quantify the navigation performance in our experiments, we posit that factors such as the distance traversed or closeness to a human-defined trajectory do not matter as much as traveling on only the preferred terrains. We therefore assign terrain types in each experiment a low, medium, and high cost, according to the human deployer’s preference and report the proportion of the planned path which belonging to each tier. The assignment of semantic terrain types to these three tiers is based on a hidden total ordering of all the terrains which appear in the environment. Each of the methods and baselines tested are deployed with a preference (provided according to their respective representations) consistent with the hidden total ordering. Note that we have purposely chosen this metric to be different from the commonly-used Hausdorff distance between the planned trajectory and one defined by a human operator, which can vary greatly when there are multiple valid paths to the goal.
Figure 4: The effect of changing preference on path planning is shown on the left image of an aerial map. The blue path corresponds to the middle costmap and the red path corresponds to the right costmap. The generated costmaps reflect the preferences provided in the context. Dark purple corresponds to low cost and yellow is high cost.